OpenStudy (anonymous):

what is the simplified form of 7x+4/x+3+6x-1/x+3

OpenStudy (anonymous):

Collect like terms. What do you get?

OpenStudy (anonymous):

for example, the 7x and 6x can be combined to give you....?

OpenStudy (anonymous):

same denominators, add up the numerators \[\frac{7x+4}{x+3}+\frac{6x-1}{x+3}=\frac{7x+4+6x-1}{x+3}\]

OpenStudy (anonymous):

then combine like terms in the numerator

OpenStudy (anonymous):

13x+3/x+3?

OpenStudy (anonymous):

that is what i get, yes

OpenStudy (anonymous):

orr... sorry, no. I got + 6, not +3.

OpenStudy (anonymous):

pretty sure it is \[+3\]

OpenStudy (anonymous):

there are 2 +3 terms